# Scaling laws for velocity profile of granular flow in rotating drums

### v2

@@ II.3.5 Numerical conditions
In this simulation, the free surface of the granular phase is initially set to be aligned with the $X$ axis, i.e., $\alpha=1$ for $Z<0$ and $\alpha=0$ for $Z\geq 0$. The initial velocity is 00 in the entire system. From this initial state, we generate a steady flow by rotating the drum counterclockwise with the angular velocity $\Omega$.
III Velocity profile
Figures 2 (a) and (b) show the velocity field of granular material for $D=150d$ and $\Omega=6.28\times 10^{-3}\sqrt{g/d}$ obtained from the DEM and CFD simulations, respectively. The red solid lines in these figures represent the free surface. In both cases, the granular material convects in a counterclockwise direction. Near the wall, the flow exhibits the rigid rotation following the rotating drum. Near the free surface, there is a region where the velocity is almost parallel to the free surface.
We set another Cartesian coordinate system $(x,z)$, where the $x$ axis is parallel to the free surface at $X=0$, as shown in Fig. 1. In this coordinate system, the depth from the free surface is represented by $z$. The coordinate system $(x,z)$ is given by
[…]
Here, $\theta$ denotes the angle of the free surface at $X=0$. The position of the free surface in the DEM simulation is determined from the packing fraction, as shown in Appendix C. Note that the free surface position at $x=0$ is treated as $z=0$, which is numerically evaluated in Appendix D.
Next, we analyze the velocity $u$ in the $x$-direction as a function of $z$ at the center of the granular bed, i.e. $x=0$. The velocity profile $u(z)$ for $D=150d$, $200d$, and $300d$ with $\Omega=6.28\times 10^{-3}\sqrt{g/d}$ is shown in Fig. 3 (a). The open and closed symbols represent the results of the DEM and CFD simulations, respectively. The results of the CFD simulation are quantitatively consistent with those of the DEM simulation. The velocity $u(z)$ decreases as $z$ increases. For larger $z$, $u(z)$ coincides with the velocity $-\Omega z$ for the rigid rotation following the rotating drum, which is shown as the solid line in Fig. 3. This region is regarded as the static flow regime. Near $z=0$, the deviation from the rigid rotation becomes significant. This region corresponds to the surface flow layer. At the boundary between these two phases, $u(z)$ is close to 00, and the two phases are smoothly connected. These results are consistent with those in Ref. [10].
In Fig. 3 (b), the velocity profile $u(z)$ obtained from the DEM and CFD simulations for various values of $\Omega$ with $D=150d$ is shown. The results of the DEM and CFD simulations are also quantitatively consistent in this figure. For each $\Omega$, the surface flow layer and the static flow regime are distinguished. The magnitude and the slope of $u(z)$ in the surface flow layer increase with $\Omega$. Figure 3 demonstrates that the continuum model using the $\mu(I)$-rheology quantitatively reproduces the velocity field over a broad parameter space, extending beyond the limited conditions examined in previous studies [23, 13, 14, 16].
[…]
Figure 2: Velocity fields obtained from the DEM (a) and CFD (b) simulations with $D=150d$ and $\Omega=1.26\times 10^{-3}\sqrt{g/d}$. The red solid line represents the free surface.
[…]
Figure 3: (a) Velocity profile $u(z)$ at $x=0$ for $\Omega=6.28\times 10^{-3}\sqrt{g/d}$ with $D=150d$, $200d$, and $300d$. (b) Velocity profile $u(z)$ at $x=0$ for $D=150d$ with $\Omega=1.26\times 10^{-3}\sqrt{g/d}$ to $1.26\times 10^{-2}\sqrt{g/d}$. The open and closed symbols represent the results of the DEM and CFD simulations, respectively. The solid lines represent $-\Omega z$ for the rigid rotation.
Next, to investigate the dependence of the velocity profile on $D$ and $\Omega$, we define the flow layer thickness $h$. Following Refs. [5, 7], we define the point with $u(z)=0$ in Fig. 3 as the boundary between the two phases and estimate the surface flow layer thickness $h$ as the distance from the free surface to the boundary point at $x=0$, which indicates
$u(z=h)=0.$ (49)
The surface flow layer thickness $h$ obtained from the DEM and CFD simulations is shown as a function of $\Omega$ for different $D$ in Fig. 4. The numerical results obtained from the DEM and CFD simulations show good agreement. The flow layer thickness $h$ increases with $D$ and depends slightly on $\Omega$. Note that the boundary has been estimated using different methods in previous studies [39, 40, 23, 10, 41]. However, the choice of the methods does not affect the qualitative behavior of $h$ as reported in Ref. [5]; this can be explained by the scaling of the velocity profile shown below (Appendix E).
[…]
IV Scaling laws
In this section, we theoretically derive scaling laws for the velocity profile of granular flow in a rotating drum using the continuum model. We consider the mass conservation equation (19) and the momentum conservation equation (20). The viscosity $\eta$ in Eq. (20) is given by Eq. (27) with Eq. (2). We neglect the first term on the left-hand side of Eq. (20) by considering the steady flow. Using the drum diameter $D$, angular velocity $\Omega$, and density $\rho_{\mathrm{g}}$, we non-dimensionalize ${\bf\it r}$, $\nabla$, ${\bf\it v}$, $\dot{\gamma}$, $p$, and $\eta_{\mathrm{g}}$ as follows:
$\tilde{{\bf\it r}}=\frac{{\bf\it r}}{D},\;\;\tilde{\nabla}=\frac{\nabla}{D},\;\;\tilde{{\bf\it v}}=\frac{{\bf\it v}}{\Omega D},\;\;\tilde{\dot{\gamma}}=\frac{\dot{\gamma}}{\Omega},\;\;\tilde{p}=\frac{p}{\rho_{\mathrm{g}}\Omega^{2}D^{2}},\;\;\tilde{\eta}_{\mathrm{g}}=\frac{\eta_{\mathrm{g}}}{\rho_{\mathrm{g}}\Omega D^{2}}$ (50)
with the dimensionless variables $\tilde{{\bf\it r}},\tilde{\nabla},\tilde{{\bf\it v}},\tilde{\dot{\gamma}},\tilde{p}$, and $\tilde{\eta}_{\mathrm{g}}$. Here, we have used $D$ as a unit of length because we focus on the overall velocity profile in the system with $D\gg d$. Substituting these equations into Eqs. (19), (20), and (27) with Eq. (2), we obtain
~$\tilde{\nabla}\cdot\tilde{{\bf\it v}}=0,$ (51)
~$-\tilde{\nabla}\cdot\left(\tilde{{\bf\it v}}\tilde{{\bf\it v}}\right)-\tilde{\nabla}\tilde{p}+\tilde{\nabla}\cdot\left[\tilde{\eta_{\mathrm{g}}}\left(\tilde{\nabla}\tilde{{\bf\it v}}+\left(\tilde{\nabla}\tilde{{\bf\it v}}\right)^{T}\right)\right]-\frac{2}{Fr}{\bf\it e}_{z}=0,$ (52)
$\tilde{\eta}_{\mathrm{g}}=\mu_{s}\frac{\tilde{p}}{\tilde{\dot{\gamma}}}+\left(\mu_{2}-\mu_{s}\right)\frac{\tilde{p}}{I_{0}\sqrt{\phi\tilde{p}}/(d/D)+\tilde{\dot{\gamma}}},$ (53)
where we have introduced the unit vector ${\bf\it e}_{z}$ in the $z$ direction and the Froude number
$Fr=\frac{\Omega^{2}D}{2g}.$ (54)
Here, we consider the dependence of the velocity profile on $D$ and $\Omega$, and $d/D$ and $Fr$ are the dimensionless parameters related to them in Eqs. (51)-(53). Thus, the dimensionless velocity field $\tilde{{\bf\it v}}({\bf\it r})$ is characterized by the Froude number $Fr$ and the diameter ratio $d/D$, except for $\mu_{s}$, $\mu_{2}$, and $I_{0}$ determined by the properties of particles; this is consistent with the experiments in Ref. [42].
Since we have studied the system for $d/D\leq 1/150\ll 1$ in the previous section, we consider the limit $d/D\to 0$ in Eqs. (51)-(53). In this limit, the only dimensionless parameter in Eqs. (51)-(53) is $Fr$. Hence, the solution $\tilde{{\bf\it v}}\left(\tilde{{\bf\it r}}\right)$ of these equation depends only on $Fr$ and is written as
where ${\bf\it V}\left(\tilde{{\bf\it r}};Fr\right)$ is a function depending on $Fr$. Taking the limit $d/D\to 0$ in Eq. (53) is equivalent to treating the granular material as a simple plastic body with the yield stress proportional to the pressure. A similar approximation has been introduced in Ref. [43]. From Eq. (55), the scaling law for the velocity profile $u(z)$ at $x=0$ is obtained as follows:
$\frac{u(z)}{\Omega D}=U\left(\frac{z}{D};Fr\right),$ (56)
where $U\left(z/D;Fr\right)$ is a scaling function.
Substituting Eq. (56) into Eq. (49), we obtain
$U\left(\frac{h}{D};Fr\right)=0.$ (57)
From this equation, we derive the scaling law for $h$ as follows:
$\frac{h}{D}=H\left(Fr\right),$ (58)
where $H(Fr)$ is a scaling function.
We verify these scaling laws from the numerical simulations. Figure 5 illustrates the normalized velocity profile $u/\Omega D$ against $z/D$ for different $D/d$ with (a) $Fr=3.00\times 10^{-3}$ and (b) $Fr=1.20\times 10^{-2}$. The closed and open symbols represent the results of the CFD and DEM simulations, respectively. The solid line corresponds to the rigid rotation. We find excellent data consolidation in Fig. 5, which verifies the validity of the scaling law, Eq. (56). Comparing the results for different $Fr$, we find that the scaling function $U\left(z/D;Fr\right)$ shows a slight dependence on $Fr$.
[…]
Next, we plot $h/D$ against $Fr$ for different $d/D$ in Fig. 6. The closed and open symbols indicate the results of the CFD and DEM simulations, respectively. This figure verifies the validity of the scaling law, Eq. (58). We have focused on the steady flow for low $Fr<10^{-2}$. In this region, the scaling function $H(Fr)$ in Eq. (58) is expanded as follows:
$H\left(Fr\right)=H^{(0)}+FrH^{(1)}+O(Fr^{2})$ (59)
where $H^{(0)}$ and $H^{(1)}$ are constants. As shown in Appendix F, $H^{(0)}$ satisfies $H^{(0)}>0$ from the conservation of the flux in the steady flow. The scaling laws (58) with Eq. (59) derived from the continuum model explain that $h/D$ is almost constant with $Fr$ for $Fr\ll 1$ as shown in Fig. 6. This scaling law indicates that the surface layer thickness $h$ is proportional to $D$ but independent of $\Omega$ for sufficiently small $Fr$. This dependence of $h$ on $D$ and $\Omega$ is consistent with the previous experiment in Ref. [2].
[…]
Figure 6: Normalized surface flow layer thickness $h/D$ against the Froude number $Fr$ for various values of $D$. The open and closed symbols represent the results of the DEM and CFD simulations, respectively. The solid line represents Eq. (59) with $H^{(0)}=1.1$ and $H^{(1)}=1.2$, which are estimated by the least squares method.
V Conclusion
In this paper, we numerically studied the steady flow of a two-dimensional granular material in a rotating drum using the DEM and CFD simulations with the $\mu(I)$-rheology. The velocity fields obtained from both methods are quantitatively consistent for various values of the diameter $D$ and the angular velocity $\Omega$ of the drum. The granular material exhibits the rigid rotation in the static flow regime near the drum wall while it flows parallel to the free surface in the surface flow layer. The surface flow layer is characterized by the layer thickness $h$ at the center of the drum. We have analytically obtained scaling laws (Eqs. (56) and (58)) for the velocity profile and surface layer thickness from dimensional considerations of the continuum equations. Numerical results verified the validity of these scaling laws.
The scaling law for the surface flow layer thickness based on the $\mu(I)$-rheology predicts $h\sim\mbox{const.}$. in the limit of $\Omega\to 0$, which is consistent with the experimental results in Ref. [2]. In contrast, other experimental studies [3, 5, 6, 7] have reported a power-law dependence, $h\sim\Omega^{\alpha}$ with a constant $\alpha$, which was obtained from a phenomenological assumption of a constant shear rate. The cylinder used in Ref. [2] is sufficiently long in the axial direction, and the flow can be considered two-dimensional, which may explain the agreement with our theoretical prediction. Our preliminary three-dimensional simulations suggest that the scaling law (Eqs. (56) and (58)) holds for granular flow in a sufficiently long rotating cylinder, whereas significant deviations are observed for shorter cylinders. These deviations may be attributed to the influence of sidewall friction, potentially explaining why the experimental results in Refs. [3, 5, 6, 7] deviate from our scaling laws. However, the effect of sidewall friction in shorter cylinders and a theoretical explanation for the experimental results in Refs. [3, 5, 6, 7] warrant further careful investigation.
We studied the rotating drum half-filled with granular materials in this study. However, a flow similar to Fig. 1 is observed for various values of the fill fraction. We have not investigated the dependence of the velocity field on the fill fraction; however, we expect the scaling laws derived in this study to apply to systems with different fill fractions because we have not used the condition for the fill fractions in the derivation of the scaling law for the velocity field ${\bf\it v}({\bf\it r})$.
Recently, the $\mu(I)$-rheology has been improved to include the non-local effect [20, 22]. We ignored its effect and adopted the local $\mu(I)$-rheology in our analysis since we deemed it irrelevant for sufficiently large systems with $D/d\geq 150$. In fact, the $\mu(I)$-rheology has qualitatively reproduced the velocity fields in our simulation. However, more detailed work is necessary on the non-local effect on the velocity field in the rotating drum.
Although we obtained the scaling laws by the non-dimensionalization, the forms of the scaling functions are not theoretically derived. In Ref. [44], the velocity field of an incompressible Newtonian fluid in a rotating drum is analytically calculated. In. Ref. [45], the streamlines for the granular flow in a rotating drum are derived based on a simple constitutive law. The development of these analytical methods may lead to a better understanding of the granular flow, including the derivation of the scaling functions.
Acknowledgements.
The authors thank S. Goto, Y. Motoori, D. Watanabe, and I. Procaccia for fruitful discussions. We would like to thank Editage (www.editage.jp) for English language editing. This study was supported by JSPS KAKENHI under Grant No. JP23K03248 and JP21H01006. H.O. conducted the simulations and analysis. All authors analyzed the results and reviewed the manuscript.
Appendix A Velocity field in DEM simulation
The velocity distribution in Fig. 2(a) is calculated using the following procedure in the DEM simulation. First, the system is divided into small square regions with a length of $4.8d$ in the plane of the Cartesian coordinates $\left(X,Z\right)$. Then, the average particle velocity in each region is plotted as the velocity field.
In the DEM simulation, the velocity field in the plane of the Cartesian coordinates $(x,z)$ shown in Fig. 3 is numerically obtained as the average particle velocity in the square regions with a length of $1.2d$. In the CFD simulation, the velocity at each region is calculated as the average of the values on the computational grids overlapping with the region.
Appendix B Determination of parameters for $\mu(I)$-rheology
To estimate the parameters in Eq. (2), we have performed the DEM simulation for the granular materials shown in Sec. II.2 under uniform shear with the shear rate $\dot{\gamma}$ and packing fraction $\phi$ using the Lees-Edwards boundary condition [46]. We use $\phi=0.825$, which is the value in the static flow regime of the rotating drum. In this simulation, the shear stress $\tau$ and pressure $p$ are measured, and we plot the bulk friction $\mu=\tau/p$ against the inertial number $I$ in Fig. 7. This figure shows that the $I$-dependence of $\mu$ is consistent with the theoretical expression, Eq. (2), for the $\mu(I)$-rheology represented by the solid line. We have determined the parameters as $\mu_{s}=0.246$, $\mu_{2}=0.401$, and $I_{0}=0.133$ by fitting the numerical data with Eq. (2).
[…]
Figure 7: Bulk friction $\mu$ against the inertial number $I$ obtained from the DEM simulation under uniform shear. The solid line represents Eq. (2) with $\mu_{s}=0.246$, $\mu_{2}=0.401$, and $I_{0}=0.133$.
Appendix C Packing fraction in the rotating drum
Figure 8 presents the dependence of the volume fraction $\phi$ on $z$ at $x=0$ obtained from the DEM simulation with $D=150d$ and $\Omega=6.28\times 10^{-3}\sqrt{\rho_{g}g/d}$. The packing fraction $\phi$ is nearly constant except for the region near $z\approx 0$. The packing fraction in the static flow regime is estimated to be $0.825$, which is indicated by the straight line in Fig. 8, though the packing fraction in the surface flow layer is slightly small due to the dilatancy of granular materials under shear [10]. Therefore, in the CFD simulation, the granular material is treated as an incompressible fluid. In the DEM simulation, $\phi=0.7$ is defined as the free surface position.
[…]
@@ Appendix C Packing fraction in the rotating drum
Appendix D Position of free surface
Figure 9 illustrates the position of the free surface $z_{\mathrm{surface}}$ against the Froude number $Fr$ at $x=0$ for each $D$. The open and closed symbols represent the results of the DEM and CFD simulations, respectively. The normalized free surface position $z_{\mathrm{surface}}/D$ is close to 00 for all of the Froude number $Fr$. Therefore, we treat $z_{\mathrm{surface}}=0$ in this study.
[…]
@@ Appendix D Position of free surface
Appendix E Boundary between surface flow layer and static flow regime
In this study, the boundary between the surface flow layer and the static flow regime is defined as the position for $u(z)=0$. However, in Ref. [39], the boundary is determined from the difference between $u(z)$ and the rigid rotation velocity $-\Omega z$. For example, we may regard the surface flow layer as the region where the velocity in the granular bed is sufficiently larger than that of the rigid rotation as
@@ Appendix E Boundary between surface flow layer and static flow regime
Appendix F Surface flow layer thickness $h$ for $Fr\ll 1$
@@
[…]
@@
[…]
@@
[…]
Since $20\leq h/D\leq 1/$ 2 from the definition, $H^{(0)}$ satisfies $20\leq H^{(0)}\leq 1/$ 2. Here, $H^{(0)}$ is the solution of Eq. (76); however, $H^{(0)}=0$ does not satisfy this equation due to Eq. (70). Therefore, we obtain
@@
This equation with Eqs(59) and (58) explains the behavior of $h$ in Fig. 6.

### v3

@@ C.5 Numerical conditions
In this simulation, the free surface of the granular phase is initially set to be aligned with the $X$ axis, i.e., $\alpha=1$ for $Z<0$ and $\alpha=0$ for $Z\geq 0$. The initial velocity is 0 in the entire system. From this initial state, we generate a steady flow by rotating the drum counterclockwise with the angular velocity $\Omega$.
Appendix D Determination of parameters for $\mu(I)$-rheology
To estimate the parameters in Eq. (4), we have performed the DEM simulation for the granular materials shown in Sec. II.1 under uniform shear with the shear rate $\dot{\gamma}$ and packing fraction $\phi$ using the Lees-Edwards boundary condition. [52] We use $\phi=0.825$, which is the value in the static flow regime of the rotating drum. In this simulation, the shear stress $\tau$ and pressure $p$ are measured, and we plot the bulk friction $\mu=\tau/p$ against the inertial number $I$ in Fig. 7. This figure shows that the $I$-dependence of $\mu$ is consistent with the theoretical expression, Eq. (4), for the $\mu(I)$-rheology represented by the solid line. We have determined the parameters as $\mu_{s}=0.246$, $\mu_{2}=0.401$, and $I_{0}=0.133$ by fitting the numerical data with Eq. (4).
Figure 7: Bulk friction $\mu$ against the inertial number II obtained from the DEM simulation under uniform shear. The solid line represents Eq. (4) with $\mu_{s}=0.246$, $\mu_{2}=0.401$, and $I_{0}=0.133$.
Appendix E Packing fraction in the rotating drum
Figure 8 presents the dependence of the volume fraction $\phi$ on $z$ at $x=0$ obtained from the DEM simulation with $D=150d$ and $\Omega=6.28\times 10^{-3}\sqrt{\rho_{g}g/d}$. The packing fraction $\phi$ is nearly constant except for the region near $z\approx 0$. The packing fraction in the static flow regime is estimated to be $0.825$, which is indicated by the straight line in Fig. 8, though the packing fraction in the surface flow layer is slightly small due to the dilatancy of granular materials under shear. [11] Therefore, in the CFD simulation, the granular material is treated as an incompressible fluid. In the DEM simulation, $\phi=0.7$ is defined as the free surface position.
@@ Appendix E Packing fraction in the rotating drum
Appendix F Position of free surface
Figure 9 illustrates the position of the free surface $z_{\mathrm{surface}}$ against the Froude number $Fr$ at $x=0$ for each $D$. The open and closed symbols represent the results of the DEM and CFD simulations, respectively. The normalized free surface position $z_{\mathrm{surface}}/D$ is close to 0 for all of the Froude number $Fr$. Therefore, we treat $z_{\mathrm{surface}}=0$ in this study.
@@ Appendix F Position of free surface
Appendix G Boundary between surface flow layer and static flow regime
In this study, the boundary between the surface flow layer and the static flow regime is defined as the position for $u(z)=0$. However, in Ref. 32, the boundary is determined from the difference between $u(z)$ and the rigid rotation velocity $\Omega z$. For example, we may regard the surface flow layer as the region where the velocity in the granular bed is sufficiently larger than that of the rigid rotation as
@@ Appendix G Boundary between surface flow layer and static flow regime
Appendix H Surface flow layer thickness $h$ for $Fr\ll 1$
@@
$Q_{+}$ $=$ $\Omega D^{2}\int_{0}^{H(Fr)}d\zeta\ U_{+}(\zeta;Fr),$ (75)
$Q_{-}$ $=$ $\Omega D^{2}\int_{0}^{\frac{1}{2}-H(Fr)}d\zeta\ U_{-}(\zeta;Fr).$ (76)
@@
$\left\{\int_{0}^{H^{(0)}}d\zeta\ U_{+}^{(0)}(\zeta)-\int_{0}^{\frac{1}{2}-H^{(0)}}d\zeta\ U_{-}^{(0)}(\zeta)\right\}+O(Fr)=0.$ (79)
@@
$\int_{0}^{H^{(0)}}d\zeta\ U_{+}^{(0)}(\zeta)=\int_{0}^{\frac{1}{2}-H^{(0)}}d\zeta\ U_{-}^{(0)}(\zeta).$ (80)
Since $0\leq h/D\leq 1/2$ from the definition, $H^{(0)}$ satisfies $0\leq H^{(0)}\leq 1/2$. Here, $H^{(0)}$ is the solution of Eq. (80); however, $H^{(0)}=0$ does not satisfy this equation due to Eq. (74). Therefore, we obtain
@@
This equation with Eqs. (20) and (19) explains the behavior of $h$ in Fig. 6.